# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

### v3

@@ Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole.. We take a string-based molecular representation — Group SELFIES — as input tokens to pre-train and fine-tune our Lamole, as it provides chemically meaningful semantics. By disentangling the information flows of Lamole, we propose considering both self-attention weights and gradients for better quantification of each chemically meaningful substructure’s impact on the model’s output. To make the explanations more faithfully respect the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction.
1 Introduction
@@
Molecular property prediction aims to reveal the molecular structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design (Fang et al., 2022; Deng et al., 2023; Tripp et al., 2023; Wang et al., 2024; Ekström Kelvinius et al., 2023; Hong et al., 2024). Several learning-based models are devised based on the underlying molecular representations, such as graph-based and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES (Weininger, 1988)), stand out for their simplicity and adaptability (Deng et al., 2023; Wigh et al., 2022; Cheng et al., 2023). By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert (Kenton & Toutanova, 2019) offer higher throughput and accuracy for molecular property prediction (Deng et al., 2023; Chithrananda et al., 2020).
@@
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX (Lin et al., 2022); (c) GNNExplainer (Ying et al., 2019), (d) GNN with gradient-based explainability technique (GradCAM (Selvaraju et al., 2017)); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES (Cheng et al., 2023) as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
@@
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design (Wu et al., 2023; Wellawatte et al., 2023; Das et al., 2022). With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction (Proietti et al., 2024; Ying et al., 2019; Ye et al., 2023; Lin et al., 2021).
However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously. Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig. 1 (b)$\sim$(e)). 2) From the perspective of explainability techniques, current methods suffer from two main limitations. First, they cannot effectively capture the interactions between functional groups within the molecular structure. Second, they could not generate explanations that align with chemists’ intuition. As a result, they fail to produce explanations that faithfully reflect the structure-property relationships. (see Fig. 1 (f)). Therefore, an effective framework is imperative for explainable molecule property predictions.
The recently proposed string-based molecular representation — Group SELFIES (Cheng et al., 2023) — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes chemically meaningful substructures as tokens, including a benzene, a nitro group, and a carboxyl group. Compared with 2D molecular graphs, Group SELFIES provides inherent semantic information, making it easier for the model to capture and understand chemically meaningful semantics. Moreover, using Group SELFIES eliminates the need to identify or segment chemically meaningful substructures in 2D molecular graphs. With Group SELFIES’s simplicity and adaptability, this work develops an explainable molecular property prediction framework based on language models to provide chemical concepts-aligned explanations, called Lamole (see Fig. 1 (g)). The contributions can be summarized as follows.
  1. 1.
    We found that existing explainability techniques fail to provide chemically meaningful explanations, perceive functional group interactions, and reveal molecular structure-property relationships faithfully. To address the issues, we use Group SELFIES to pre-train and fine-tune LMs to make LMs easily understand the chemically meaningful semantics. Moreover, the process of generating explanations should reflect the reasoning process behind the model architecture. Therefore, by disentangling the information flows of Transformer-based LMs, we integrate the self-attention weights and gradients to capture the substructure interactions to better quantify each chemically meaningful substructure’s contribution to the predicted molecular properties.
  2. 2.
    To make the explanations more faithfully respect the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations. We show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
  3. 3.
    We first bridge the manifold hypothesis with explainable molecular property prediction. We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.3\%$. We also quantitatively evaluate explanations based on the first proposed plausibility metric. Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$. Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
2 Related Work
Several explainable GNNs are proposed to explain the relationship between the input graph and the prediction (Sun et al., 2023; Xiong et al., 2019; Lin et al., 2022; Ying et al., 2019; Luo et al., 2020; Lin et al., 2021). Among these works, structure similarity or attention weights are proposed to capture structural interaction (Sun et al., 2023; Xiong et al., 2019). However, two similar substructures do not necessarily lead to interaction between them, and attention weights are often inconsistent with the feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020). In addition, as shown in Fig. 1 (b)$\sim$(d), some trivial structures received relatively high importance scores, indicating the explanations might not align well with the chemical concepts.
@@ 2 Related Work
On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction (Chithrananda et al., 2020; Wang et al., 2019; Ahmad et al., 2022; Ross et al., 2022). However, the "black-box" characteristics of LMs hamper trust use of these potent computational tools in scientific domains. Some explainability techniques could be applied to LMs. One way is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020). Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the input importance. However, the generated explanations may change drastically with very small perturbations (Agarwal et al., 2021). Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature (Selvaraju et al., 2017). However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations (Adebayo et al., 2018; Agarwal et al., 2021; Rudin, 2019). Therefore, the explanation generation process should reflect the model reasoning process behind the model architectures. To this end, we disentangle the model architectures’ information flows to generate explanations that faithfully reveal the structure-property relationship.
3 Methodology
@@
Problem Setup. Given a dataset $\mathcal{G}=\{(g^{(i)},y^{(i)})\}$ consisting of molecular graphs $\{g^{(i)}\}$ with their property labels $\{y^{(i)}\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as denoted as $f:g\mapsto y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
@@ 3.1 Our Design: Lamole
Fine-Tuning Stage. Fig. 3 shows the fine-tuning stage of the proposed Lamole. We assume that the Transformer encoder in Lamole stacks $L$ identical Transformer layers to encode the molecular string $s$ as token embedding $\mathbf{h}^{(l)}=\{\mathbf{h}_{1}^{(l)},\ldots,\mathbf{h}_{j}^{(l)}\}$, where $\mathbf{h}^{(l)}$ is the token embedding at the $l$-th layer. We use the self-attention weighted average embedding $\mathbf{h}_{o}=(\sum_{j=1}\alpha_{j}\cdot\mathbf{h}_{j}^{(L)})/(\sum_{j=1}\alpha_{j})$ for molecular property prediction, where $\alpha_{j}$ is the attention weight of the $j$-th token. A multilayer perceptron (MLP) classifier is added to predict the molecular property $\hat{y}=\mathrm{MLP}(\mathbf{h}_{o})$ by minimizing the classification cross-entropy loss $\mathcal{L}_{CE}(y,\hat{y})$.
Information Flow-Based Explanations. To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature (Selvaraju et al., 2017). However, we argue that using gradients alone cannot effectively capture substructure interactions. As depicted in Fig. 1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
We illustrate the possible reason by disentangling the information flows of the Transformer. As shown in Fig. 3 right panel, due to the skip connection and attention mechanism, both attention weights, gradients, and the input contribute to the outputs. Therefore, using gradients alone as explanations could fall short of capturing interactions. To address the issue, we leverage both attention and gradients, as well as the input, to derive explanations. Below, we will elaborate on integrating attention weights into gradient-based explanations.
@@
Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM (Selvaraju et al., 2017), the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input, and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
@@
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans (Wiegreffe & Pinter, 2019; Herman, 2017; Jacovi & Goldberg, 2020). Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
@@
Eq. (4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth. The lower the scores on non-ground truth and the greater the scores on ground truth, the better explanation plausibility. Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
@@
$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\left[\max\left(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}-\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\right)+\triangle_{1}\right],$ (5)

### v4

@@ Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations. We take a string-based molecular representation — Group SELFIES — as input tokens to pre-train and fine-tune our Lamole, as it provides chemically meaningful semantics. By disentangling the information flows of Lamole, we propose considering both self-attention weights and gradients for better quantification of each chemically meaningful substructure’s impact on the model’s output. To make the explanations more faithful to the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over eight datasets demonstrate Lamole can achieve comparable prediction accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction. To further illustrate the actionable utility of the explanations derived from Lamole, we integrated the framework with an evolutionary algorithm. This integration established an interpretable optimization pipeline for molecular editing, demonstrating that Lamole functions beyond simple post-hoc analysis but serves as a practical guide for molecule discovery.
I Introduction
@@
Molecular property prediction aims to reveal the structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design [53, 12, 10, 51, 11, 17, 24, 26, 27, 25]. Several learning-based models are devised based on the underlying molecular representations, such as graph-based and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [55]), stand out for their simplicity and adaptability [10, 58, 6]. By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [22] offer higher throughput and accuracy for molecular property prediction [10, 7].
@@
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX [30]; (c) GNNExplainer [64], (d) GNN with gradient-based explainability technique (GradCAM [44]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [6] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [60, 56, 8]. With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [39, 64, 63, 29].
[…]
The recently proposed string-based molecular representation — Group SELFIES [6] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes chemically meaningful substructures as tokens, including a benzene, a nitro group, and a carboxyl group. Compared with 2D molecular graphs, Group SELFIES provides inherent semantic information, making it easier for the model to capture and understand chemically meaningful semantics. Moreover, using Group SELFIES eliminates the need to identify or segment chemically meaningful substructures in 2D molecular graphs. With Group SELFIES’s simplicity and adaptability, this work develops an explainable molecular property prediction framework based on language models to provide chemical concepts-aligned explanations, called Lamole (see Fig. 1 (g)). The contributions can be summarized as follows.
  1. 1.
    We found that existing explainability techniques fail to reveal molecular structure-property relationships faithfully. Therefore, to provide explanations directly aligned with chemical intuition, we first leverage the Group SELFIES representation—which inherently encodes chemically meaningful substructures—for both pre-training and fine-tuning our LMs. By integrating this chemically grounded semantic input, our framework enables LMs to directly comprehend and operate on functional group semantics, moving beyond low-level features to generate explanations that are immediately relevant and intuitive for chemists.
  2. 2.
    We analyzed the limitations of existing methods in revealing molecular structure-property relationships, noting that approaches relying solely on attention or gradients as explanations often fail to adequately reflect the model’s underlying reasoning process. For an explanation to be faithful, the generation process must be closely coupled with the mechanism of the model architecture. Therefore, by carefully disentangling the information flows within Transformer-based LMs, we propose an information flow-based explainable method. This strategy is designed specifically to capture the interactions between chemically meaningful substructures, thereby providing a more accurate quantification of each substructure’s contribution to the predicted molecular properties.
  3. 3.
    To make the explanations more faithful to the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations. Empirically, we show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%. Furthermore, theoretically, we first bridge the manifold hypothesis with explainable molecular property prediction. We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
  4. 4.
    We further demonstrate the actionable utility of the explanations derived from Lamole through a dedicated molecular editing case study. By integrating Lamole’s identified key functional groups with an evolutionary algorithm, we establish an interpretable optimization pipeline that guides the molecular optimization process to enhance target properties. This integration showcases Lamole’s capability to bridge predictive insights with rational molecular design and optimization, paving the way for more efficient chemical discovery.
  5. 5.
    We evaluated our method across eight diverse datasets spanning both classification and regression tasks, including a wide range of molecular properties such as hepatotoxicity, mutagenicity, water solubility, and permeability. Experimental results demonstrate that our proposed Lamole achieves comparable predictive accuracy relative to other established baselines while exhibiting superior explanation accuracy, confirming the broad applicability of Lamole across different property prediction tasks.
@@
II Related Work
Molecular property prediction is a fundamental task aimed at elucidating the relationship between molecular structure and function. Various learning-based models have been developed based on distinct molecular representations, primarily categorized into graph-based and string-based formats. GNNs explicitly model molecular systems by incorporating atomic positions and pairwise interactions through message passing. For instance, SchNOrb [43] adapted distance-based convolution by integrating symmetry-adapted atomic orbital features for electronic structure prediction. MiCaM [13] introduced a fragment-based deep generative model that enhances novelty by automatically discovering connection-aware motifs through iterative subgraph merging, which are then utilized in a dedicated generator for de novo molecular generation. Alternatively, string-based molecular representations, such as SMILES, are favored for their simplicity and adaptability within sequential models. TamGen [59], for example, leveraged a GPT-like chemical language model to achieve target-aware molecule generation and compound refinement. This work demonstrated the efficacy of GPT-like architectures and training strategies for generation within the one-dimensional SMILES chemical space by integrating specific modules for encoding protein structure and seeding compound information.
[…]
@@ II Related Work
On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [7, 52, 4, 40]. However, the "black-box" characteristics of LMs hamper trust use of these potent computational tools in scientific domains. Some explainability techniques could be applied to LMs. One way is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance [20, 45, 1]. Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the input importance. However, the generated explanations may change drastically with very small perturbations [3]. Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [44]. However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [2, 3, 41]. Therefore, the explanation generation process should reflect the model reasoning process behind the model architectures. To this end, we disentangle the model architectures’ information flows to generate explanations that faithfully reveal the structure-property relationship.
III Methodology
@@
Problem Setup. Given a dataset $\mathcal{G}=\{(g^{(i)},y^{(i)})\}$ consisting of molecular graphs $\{g^{(i)}\}$ with their property labels $\{y^{(i)}\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as denoted as $f:g\mapsto y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., $\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
@@ III-A Our Design: Lamole
[…]
Analysis of Foundational Explanation Methods. Before introducing our information flow-based approach, we first analyze two foundational methods—attention and gradients—and detail their limitations when applied to Transformer-based LMs.
1) Attention as Explanations: This approach quantifies the importance of each molecule token to the LM’s final decision. This is achieved by extracting the softmax cross-attention from the starting answer token to all input tokens across all layers of the backbone LM. The resulting prediction-to-token attention is represented as $A_{\text{st}}(s)\in\mathbb{R}^{L\times H\times 1\times T}$, where $L$ and $H$ denote the number of layers and heads-per-layer, respectively, and $T$ is the total number of tokens provided. We subsequently compute the prediction-to-token attention $\hat{A}_{\text{st}}(s)$ by averaging across attention heads: $\hat{A}_{\text{st}}(s)=\frac{1}{H}\sum_{h=1}^{H}A_{\text{st}}(s)$. However, recent studies suggest that “attention is not explanation" because the attention scores often fail to reflect the true importance of the input features to the model’s decision [20, 45, 1].
2). Gradients as Explanations: This common practice utilizes gradient-based methods to determine feature importance by analyzing the partial derivatives of the output with respect to each input feature [44]. We seek to identify relevant tokens by computing the gradient of the LM’s decision with respect to the input tokens. The pure gradient is versatile as it is not restricted to a Transformer architecture. Specifically, for a molecule-label pair $(s,y)$, the importance is computed as $G(s,y)=\|\nabla_{s}v(s,y)\|_{2}$, where $v(s,y)$ is the logarithm of the maximum output probability of the LM.
While both practices are widely used to derive explanations, we argue that relying on gradients alone is insufficient for effectively capturing substructure interactions. As illustrated in Fig. 1 (f), a pure gradient-based method like GradCAM can inaccurately attribute a property to an unrelated structural fragment, such as incorrectly highlighting a distant benzene ring alongside the true nitro group contributor. The primary reason for this deficiency lies in the internal architecture of the Transformer. As shown in the right panel of Fig. 3, the output is influenced by a confluence of factors, including attention weights, gradients, and the original input itself, primarily due to the ubiquitous skip connections and the attention mechanism. Consequently, using gradients in isolation fails to capture the rich network of information exchange—the interactions—that occurs between substructures via the attention layers. To address this limitation, we leverage both attention and gradients, integrating them alongside the input to derive a more comprehensive explanation. The following section details this integration of attention weights into our gradient-based explanation framework.
@@
Information Flow-Based Explanations. Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM [44], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input, and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
@@
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans [57, 16, 19]. Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
Eq. (4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth. The lower the scores on non-ground truth and the greater the scores on ground truth, the better the explanation plausibility. Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
@@
$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\Bigg[\max\Bigg(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}\\ -\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\Bigg)+\triangle_{1}\Bigg],$ (5)